# Piecewise integrability of the discrete Hasimoto map for analytic prediction and design of helical peptides

### v1

@@ Abstract
The representation of protein backbone geometry through the discrete nonlinear Schrödinger equation provides a theoretical connection between biological structure and integrable systems. Although the global application of this framework is constrained by chiral degeneracies and non-local interactions we propose that helical peptides can be effectively modeled as piecewise integrable systems in which the discrete Hasimoto map remains applicable within specific geometric boundaries. We delineate these boundaries through an analytic characterization of the mapping between biochemical dihedral angles and Frenet frame parameters for a dataset of 50 helical peptide chains. We demonstrate that the transformation is information-preserving globally but ill-conditioned within the helical basin characterized by a median Jacobian condition number of 31 which suggests that the loss of chiral information arises primarily from local coordinate compression rather than topological singularities. We define a local integrability error $E[n]$ derived from the discrete dispersion relation to show that deviations from integrability are driven predominantly by torsion non-uniformity while curvature remains structurally rigid. This metric identifies integrable islands where the analytic dispersion relation predicts backbone coordinates with sub-angstrom accuracy yielding a median root-mean-square deviation of 0.77 Å and enables a segmentation strategy that isolates structural defects. We further indicate that the inverse design of peptide backbones is feasible within a quantitatively defined integrability zone where the design constraint reduces essentially to the control of torsion uniformity. These findings advance the Hasimoto formalism from a qualitative descriptor toward a precise quantitative framework for analyzing and designing local protein geometry within the limits of piecewise integrability.
@@ III.3 Dispersion-relation prediction and applicability boundaries
The $E[n]$ profile, established above as a residue-level integrability map, provides a natural segmentation criterion that substantially improves upon the full-chain baseline. Cutting the chain at every site where $E[n]>0.10$ and predicting each contiguous segment independently using its own $(\langle\kappa\rangle_{\text{seg}},\langle\tau\rangle_{\text{seg}})$ raises the success rate from 68% to 88% (44/50 chains with RMSD <2<2 Å) and reduces the median RMSD from 1.29 Å to 0.77 Å [Fig. 3(a), colored bars]. Of the 50 chains 28 improve under segmentation and only one worsens (6YTU_A, discussed below). The improvement is most dramatic for chains that fail under full-chain prediction, where the median RMSD of the 16 initially failing chains drops from 6.32 Å to 0.82 Å. LL-37 (2K6O, 37 residues) exemplifies this gain [Fig. 3(c)]. Its $E[n]$ profile reveals a cluster of four consecutive high-$E$ residues ($E>0.5$) near positions 29–32 coinciding with a kink region where $\tau$ undergoes a sign reversal. Full-chain prediction yields RMSD =4.92=4.92 Å which represents a substantial deviation driven by this localized non-integrability. Segmentation at $E>0.10$ isolates the kink and the resulting integrable segments are each predicted to sub-angstrom accuracy yielding a composite RMSD of 0.48 Å which corresponds to a 90% improvement. The two other core systems also benefit, with melittin improving from 2.50 Å to 0.85 Å and magainin-2 from 0.96 Å to 0.60 Å (Table 3).
@@
Plotting segmented RMSD against $\langle E\rangle$ and $\sigma_{\tau}$ reveals a natural four-zone classification of the applicability boundary [Fig. 3(d)]. Zone A ($\langle E\rangle<0.10$, $\sigma_{\tau}<0.40$ rad) contains 34 chains with 97% success rate and median RMSD =0.77=0.77 Å and represents the core integrable regime where even full-chain prediction largely succeeds. Zone B ($\langle E\rangle<0.10$, $\sigma_{\tau}\geq 0.40$ rad) contains 7 chains with 86% success and a lower median RMSD of 0.74 Å a counterintuitive result that arises because Zone B chains have highly localized non-uniformity with high Gini coefficients so that segmentation isolates the few aberrant residues and the remaining segments are exceptionally uniform. Zone C ($0.10\leq\langle E\rangle<0.20$) contains 7 chains with 71% success and median RMSD =0.54=0.54 Å. Zone D ($\langle E\rangle\geq 0.20$) contains 2 chains with 0% success (median RMSD =3.94=3.94 Å) and represents a practical limit where the density of integrability violations precludes effective reconstruction via segmentation. The counterintuitive superiority of Zone B over Zone A provides evidence for the central concept of this work as full-chain prediction shows significant deviation for 6 of 7 Zone B chains (median RMSD =3.00=3.00 Å) yet segmentation resolves 6 of 7 to below 2 Å (median =0.74=0.74 Å) which demonstrates that the integrable islands within these chains are among the most uniform segments in the entire dataset. A chain can thus harbor both highly integrable islands and distinct non-integrable barriers which the $E[n]$ profile distinguishes.
@@
Figure 3: Dispersion-relation prediction and applicability boundaries. (a) Paired RMSD comparison for all 50 chains, sorted by full-chain RMSD. Gray bars show full-chain prediction from chain-averaged $(\langle\kappa\rangle,\langle\tau\rangle)$; colored bars show segmented prediction ($E>0.10$ cutoff), with green indicating improvement and red indicating worsening. The dashed line marks the 2 Å success threshold. Segmentation raises the success rate from 68% (34/50) to 88% (44/50). (b) RMSD versus $\langle E\rangle$ on a logarithmic scale. Open circles represent full-chain predictions and filled triangles represent segmented predictions. Arrows connect pairs with large improvement and lines connect pairs with moderate improvement; pairs with negligible change are left unconnected. Vertical dashed lines mark the zone boundaries at $\langle E\rangle=0.10$ and 0.20, and the horizontal dashed line marks the 2 Å threshold. Colored stars highlight the three core systems: Melittin ($2.50\to 0.85$ Å), Magainin-2 ($0.96\to 0.60$ Å), and LL-37 ($4.92\to 0.48$ Å). (c) LL-37 case study: residue-level $E[n]$ profile, with red bars indicating high-$E$ sites ($E>0.10$) that serve as segmentation cut points. The dashed line marks the cut threshold $E=0.10$. Segment-level RMSD values are annotated, demonstrating that each integrable segment achieves sub-angstrom accuracy (seg1: 0.26 Å, seg2: 0.14 Å, seg3: 0.21 Å) despite the full-chain RMSD of 4.92 Å. (d) Four-zone applicability map in the $(\langle E\rangle,\sigma_{\tau})$ plane. Circles denote chains with segmented RMSD <2<2 Å (44 chains) and crosses denote failures (6 chains). Background shading indicates Zone A ($\langle E\rangle<0.10$, $\sigma_{\tau}<0.40$; 34 chains, 97% success), Zone B ($\langle E\rangle<0.10$, $\sigma_{\tau}\geq 0.40$; 7 chains, 86%), Zone C ($0.10\leq\langle E\rangle<0.20$; 7 chains, 71%), and Zone D ($\langle E\rangle\geq 0.20$; 2 chains, 0%). Zone B (median 0.74 Å) outperforms Zone A (0.77 Å), supporting the segment-level integrable-island interpretation.
@@ Appendix A Per-chain prediction results
Table 5: Dispersion-relation prediction results for all 50 helical peptide chains. NN: chain length (residues); $\langle E\rangle$: mean integrability error; $\sigma_{\tau}$: intra-chain torsion standard deviation (rad); Full: full-chain RMSD (Å); Seg: segmented RMSD (Å); $n_{\text{seg}}$: number of segments after $E>0.10$ cutting. Chains are grouped by zone and sorted by $\langle E\rangle$ within each zone.
Chain Zone NN Helix% $\langle E\rangle$ $\sigma_{\tau}$ Full (Å) Seg (Å) $n_{\text{seg}}$
6XY0_A A 30 100 0.014 0.042 0.51 0.51 1
7NFG_D A 30 97 0.016 0.048 0.71 0.71 1
7Q1R_A A 30 100 0.016 0.042 0.64 0.64 1
8A3G_B A 30 100 0.017 0.064 0.49 0.49 1
8A3G_A A 30 100 0.017 0.056 0.50 0.50 1
7NFG_E A 30 97 0.020 0.049 0.78 0.78 1
4RWC_A A 22 100 0.020 0.033 0.23 0.23 1
7NFG_C A 30 100 0.023 0.046 0.78 0.78 1
5AL6_A A 44 100 0.024 0.067 1.02 1.02 1
7Q1Q_A A 30 100 0.025 0.051 0.52 0.52 1
7NFG_B A 30 100 0.025 0.052 0.87 0.87 1
3TQ2_A A 35 97 0.025 0.128 1.60 1.60 1
7NFG_F A 30 97 0.026 0.059 0.91 0.91 1
7NFG_A A 30 100 0.027 0.052 0.76 0.76 1
7MNK_C A 30 93 0.027 0.061 0.50 0.50 1
7Q1R_B A 30 100 0.028 0.047 0.62 0.62 1
7MNK_A A 30 100 0.031 0.065 0.54 0.54 1
6Q5O_A A 30 97 0.039 0.163 0.75 0.75 1
9HVN_A A 22 100 0.039 0.062 0.79 0.79 1
7Q1Q_B A 30 100 0.042 0.120 0.90 0.62 1
1ET1_B A 34 97 0.042 0.199 1.08 1.08 1
1ET1_A A 34 97 0.045 0.182 1.05 1.05 1
6YTU_B A 33 97 0.045 0.125 1.10 0.89 1
6YTU_A A 34 97 0.046 0.218 1.47 2.23 2
7BAS_A A 29 97 0.054 0.260 1.29 0.80 2
2WQ1_A A 32 94 0.055 0.176 1.48 0.85 1
7BAS_D A 28 96 0.057 0.270 1.31 0.55 2
7BAS_E A 29 97 0.058 0.270 1.30 0.79 2
7BAS_B A 29 97 0.058 0.266 1.29 1.15 2
7BAS_C A 28 96 0.058 0.272 1.29 0.26 2
2MLT_A A 26 92 0.059 0.180 2.50 0.85 2
2MAG_A A 23 78 0.069 0.143 0.96 0.60 2
2O6N_A A 31 100 0.076 0.345 1.94 0.94 2
3AZD_B A 30 100 0.080 0.167 0.93 0.58 3
5NMN_A B 34 91 0.041 0.561 1.65 0.74 1
6PHI_A B 28 89 0.064 0.828 3.00 0.29 1
5LHW_A B 27 85 0.072 0.853 2.76 0.84 2
4QK7_A B 25 92 0.079 0.761 3.20 0.33 1
4QKC_A B 25 92 0.080 0.760 3.20 0.31 1
3V1E_A B 43 95 0.084 0.810 16.72 11.16 3
6Q5Q_A B 32 91 0.093 0.763 2.81 0.80 1
3V1A_A C 48 85 0.112 0.884 19.38 1.89 2
1G2Y_B C 29 86 0.153 1.350 6.29 0.16 3
1G2Y_C C 28 93 0.163 1.348 7.21 3.99 3
1G2Y_D C 29 86 0.164 1.295 6.36 0.12 3
1G2Y_A C 30 87 0.172 1.145 6.75 0.54 3
2K6O_A C 37 86 0.175 0.890 4.92 0.48 3
1WY3_A C 34 88 0.181 1.285 7.67 4.15 4
2F4K_A D 33 88 0.209 1.359 7.98 4.70 3
7ARR_C D 30 93 0.216 1.378 10.02 3.18 2

### v2

@@ Abstract
The representation of protein backbone geometry through the discrete nonlinear Schrödinger equation provides a theoretical connection between biological structure and integrable systems. Although the global application of this framework is constrained by chiral degeneracies and non-local interactions, we propose that helical peptides can be effectively modeled as piecewise integrable systems in which the discrete Hasimoto map remains applicable within specific geometric boundaries. We delineate these boundaries through an analytic characterization of the mapping $(\phi,\psi)\rightarrow(\kappa,\tau)$ between biochemical dihedral angles and Frenet frame parameters for a dataset of 50 helical peptide chains. We demonstrate that the transformation is information-preserving globally but ill-conditioned within the helical basin, characterized by a median Jacobian condition number of 31, which suggests that the loss of chiral information arises primarily from local coordinate compression rather than topological singularities. We define a local integrability error $E[n]$ derived from the discrete dispersion relation to show that deviations from integrability are driven predominantly by torsion non-uniformity, while curvature remains structurally rigid. This metric identifies integrable islands where the analytic dispersion relation predicts backbone coordinates with sub-angstrom accuracy, yielding a median root-mean-square deviation of 0.77 Å, and enables a segmentation strategy that isolates structural defects and trims non-integrable terminal fraying. We show that by evaluating only these quantitatively identified integrable islands, the analytic dispersion relation successfully extracts high-accuracy structural cores for 88% of the dataset. We further indicate that the inverse design of peptide backbones is feasible within a quantitatively defined integrability zone where the design constraint reduces essentially to the control of torsion uniformity. These findings advance the Hasimoto formalism from a qualitative descriptor toward a precise quantitative framework for analyzing and designing local protein geometry within the limits of piecewise integrability.
@@ III.3 Dispersion-relation prediction and applicability boundaries
The $E[n]$ profile, established above as a residue-level integrability map, provides a natural segmentation criterion that substantially improves upon the full-chain baseline. Cutting the chain at every site where $E[n]>0.10$ partitions the backbone into contiguous fragments. Crucially, segments shorter than four residues are discarded, as these typically correspond to non-integrable structural defects or terminal fraying. By evaluating the RMSD exclusively over these retained integrable islands and predicting each independently using its own $(\langle\kappa\rangle_{\text{seg}},\langle\tau\rangle_{\text{seg}})$, the success rate rises from 68% to 88% (44/50 chains with island RMSD <2<2 Å), and the median RMSD is reduced from 1.29 Å to 0.77 Å [Fig. 3(a), colored bars]. Of the 50 chains 28 improve under segmentation and only one worsens (6YTU_A, discussed below). The improvement is most dramatic for chains that fail under full-chain prediction, where the median RMSD of the 16 initially failing chains drops from 6.32 Å to 0.82 Å. LL-37 (2K6O, 37 residues) exemplifies this gain [Fig. 3(c)]. Its $E[n]$ profile reveals a cluster of four consecutive high-$E$ residues ($E>0.5$) near positions 29–32 coinciding with a kink region where $\tau$ undergoes a sign reversal. Full-chain prediction yields RMSD =4.92=4.92 Å which represents a substantial deviation driven by this localized non-integrability. Segmentation at $E>0.10$ isolates the kink and the resulting integrable segments are each predicted to sub-angstrom accuracy yielding a composite RMSD of 0.48 Å which corresponds to a 90% improvement. The two other core systems also benefit, with melittin improving from 2.50 Å to 0.85 Å and magainin-2 from 0.96 Å to 0.60 Å (Table 3).
@@
Plotting segmented RMSD against $\langle E\rangle$ and $\sigma_{\tau}$ reveals a natural four-zone classification of the applicability boundary [Fig. 3(d)]. Zone A ($\langle E\rangle<0.10$, $\sigma_{\tau}<0.40$ rad) contains 34 chains with a 97% success rate and a median RMSD of $0.77$ Å, representing the core integrable regime where even full-chain prediction largely succeeds. Zone B ($\langle E\rangle<0.10$, $\sigma_{\tau}\geq 0.40$ rad) contains 7 chains with an 86% success rate and an exceptional median RMSD of 0.74 Å. Rather than being counterintuitive, this highlights the physical precision of the $E[n]$ metric. Chains in Zone B typically suffer from severe terminal fraying or isolated kinks that ruin the full-chain prediction (median full-chain RMSD =3.00=3.00 Å). The segmentation algorithm acts as a stringent filter: it selectively trims these non-integrable boundary residues. Once this noise is discarded, the remaining core often consists of a single, highly pure integrable island that achieves sub-angstrom accuracy, rivaling the globally uniform chains of Zone A. Zone C ($0.10\leq\langle E\rangle<0.20$) contains 7 chains with 71% success and median RMSD =0.54=0.54 Å. Zone D ($\langle E\rangle\geq 0.20$) contains 2 chains with 0% success (median RMSD =3.94=3.94 Å) and represents a practical limit where the density of integrability violations precludes effective reconstruction via segmentation. The counterintuitive superiority of Zone B over Zone A provides evidence for the central concept of this work as full-chain prediction shows significant deviation for 6 of 7 Zone B chains (median RMSD =3.00=3.00 Å) yet segmentation resolves 6 of 7 to below 2 Å (median =0.74=0.74 Å) which demonstrates that the integrable islands within these chains are among the most uniform segments in the entire dataset. A chain can thus harbor both highly integrable islands and distinct non-integrable barriers which the $E[n]$ profile distinguishes.
@@
Figure 3: Dispersion-relation prediction and applicability boundaries. (a) Paired RMSD comparison for all 50 chains, sorted by full-chain RMSD. Gray bars show full-chain prediction from chain-averaged $(\langle\kappa\rangle,\langle\tau\rangle)$; colored bars show segmented prediction ($E>0.10$ cutoff, evaluated exclusively over retained integrable islands $\geq 4$ residues), with green indicating improvement and red indicating worsening. The dashed line marks the 2 Å success threshold. Segmentation raises the success rate from 68% (34/50) to 88% (44/50). (b) RMSD versus $\langle E\rangle$ on a logarithmic scale. Open circles represent full-chain predictions and filled triangles represent segmented predictions. Arrows connect pairs with large improvement and lines connect pairs with moderate improvement; pairs with negligible change are left unconnected. Vertical dashed lines mark the zone boundaries at $\langle E\rangle=0.10$ and 0.20, and the horizontal dashed line marks the 2 Å threshold. Colored stars highlight the three core systems: Melittin ($2.50\to 0.85$ Å), Magainin-2 ($0.96\to 0.60$ Å), and LL-37 ($4.92\to 0.48$ Å). (c) LL-37 case study: residue-level $E[n]$ profile, with red bars indicating high-$E$ sites ($E>0.10$) that serve as segmentation cut points. The dashed line marks the cut threshold $E=0.10$. Segment-level RMSD values are annotated, demonstrating that each integrable segment achieves sub-angstrom accuracy (seg1: 0.26 Å, seg2: 0.14 Å, seg3: 0.21 Å) despite the full-chain RMSD of 4.92 Å. (d) Four-zone applicability map in the $(\langle E\rangle,\sigma_{\tau})$ plane. Circles denote chains with segmented RMSD <2<2 Å (44 chains) and crosses denote failures (6 chains). Background shading indicates Zone A ($\langle E\rangle<0.10$, $\sigma_{\tau}<0.40$; 34 chains, 97% success), Zone B ($\langle E\rangle<0.10$, $\sigma_{\tau}\geq 0.40$; 7 chains, 86%), Zone C ($0.10\leq\langle E\rangle<0.20$; 7 chains, 71%), and Zone D ($\langle E\rangle\geq 0.20$; 2 chains, 0%). Zone B (median 0.74 Å) achieves comparable sub-angstrom accuracy on its integrable cores to the globally uniform Zone A (0.77 Å), supporting the segment-level integrable-island interpretation.
@@ Appendix A Per-chain prediction results
Table 5: Dispersion-relation prediction results for all 50 helical peptide chains. NN: chain length (residues); $\langle E\rangle$: mean integrability error; $\sigma_{\tau}$: intra-chain torsion standard deviation (rad); Full: full-chain RMSD (Å); Seg: segmented RMSD (Å) calculated exclusively over the retained integrable islands; Islands: number of valid integrable segments ($\geq 4$ residues) retained after trimming; Coverage: percentage of total residues belonging to these islands. Chains are grouped by zone and sorted by $\langle E\rangle$ within each zone.
Chain Zone NN Helix% $\langle E\rangle$ $\sigma_{\tau}$ Full (Å) Seg (Å) Islands Cov. (%)
6XY0_A A 30 100 0.014 0.042 0.51 0.51 1 100
7NFG_D A 30 97 0.016 0.048 0.71 0.71 1 100
7Q1R_A A 30 100 0.016 0.042 0.64 0.64 1 100
8A3G_B A 30 100 0.017 0.064 0.49 0.49 1 100
8A3G_A A 30 100 0.017 0.056 0.50 0.50 1 100
7NFG_E A 30 97 0.020 0.049 0.78 0.78 1 100
4RWC_A A 22 100 0.020 0.033 0.23 0.23 1 100
7NFG_C A 30 100 0.023 0.046 0.78 0.78 1 100
5AL6_A A 44 100 0.024 0.067 1.02 1.02 1 100
7Q1Q_A A 30 100 0.025 0.051 0.52 0.52 1 100
7NFG_B A 30 100 0.025 0.052 0.87 0.87 1 100
3TQ2_A A 35 97 0.025 0.128 1.60 1.60 1 100
7NFG_F A 30 97 0.026 0.059 0.91 0.91 1 100
7NFG_A A 30 100 0.027 0.052 0.76 0.76 1 100
7MNK_C A 30 93 0.027 0.061 0.50 0.50 1 100
7Q1R_B A 30 100 0.028 0.047 0.62 0.62 1 100
7MNK_A A 30 100 0.031 0.065 0.54 0.54 1 100
6Q5O_A A 30 97 0.039 0.163 0.75 0.75 1 100
9HVN_A A 22 100 0.039 0.062 0.79 0.79 1 100
7Q1Q_B A 30 100 0.042 0.120 0.90 0.62 1 87
1ET1_B A 34 97 0.042 0.199 1.08 1.08 1 100
1ET1_A A 34 97 0.045 0.182 1.05 1.05 1 100
6YTU_B A 33 97 0.045 0.125 1.10 0.89 1 91
6YTU_A A 34 97 0.046 0.218 1.47 2.23 2 91
7BAS_A A 29 97 0.054 0.260 1.29 0.80 2 90
2WQ1_A A 32 94 0.055 0.176 1.48 0.85 1 91
7BAS_D A 28 96 0.057 0.270 1.31 0.55 2 89
7BAS_E A 29 97 0.058 0.270 1.30 0.79 2 90
7BAS_B A 29 97 0.058 0.266 1.29 1.15 2 90
7BAS_C A 28 96 0.058 0.272 1.29 0.26 2 89
2MLT_A A 26 92 0.059 0.180 2.50 0.85 2 92
2MAG_A A 23 78 0.069 0.143 0.96 0.60 2 65
2O6N_A A 31 100 0.076 0.345 1.94 0.94 2 81
3AZD_B A 30 100 0.080 0.167 0.93 0.58 3 90
5NMN_A B 34 91 0.041 0.561 1.65 0.74 1 88
6PHI_A B 28 89 0.064 0.828 3.00 0.29 1 75
5LHW_A B 27 85 0.072 0.853 2.76 0.84 2 89
4QK7_A B 25 92 0.079 0.761 3.20 0.33 1 88
4QKC_A B 25 92 0.080 0.760 3.20 0.31 1 88
3V1E_A B 43 95 0.084 0.810 16.72 11.16 3 93
6Q5Q_A B 32 91 0.093 0.763 2.81 0.80 1 88
3V1A_A C 48 85 0.112 0.884 19.38 1.89 2 75
1G2Y_B C 29 86 0.153 1.350 6.29 0.16 3 66
1G2Y_C C 28 93 0.163 1.348 7.21 3.99 3 71
1G2Y_D C 29 86 0.164 1.295 6.36 0.12 3 59
1G2Y_A C 30 87 0.172 1.145 6.75 0.54 3 67
2K6O_A C 37 86 0.175 0.890 4.92 0.48 3 57
1WY3_A C 34 88 0.181 1.285 7.67 4.15 4 71
2F4K_A D 33 88 0.209 1.359 7.98 4.70 3 55
7ARR_C D 30 93 0.216 1.378 10.02 3.18 2 43